Find the sections of the conicoid
x^2/36 + y^2/9 - z^2/4 = x^2-1 by the coordinate and identify them.
1
Expert's answer
2020-11-13T11:28:00-0500

Conicoid equation is:"\dfrac{x^2}{6^2}+\dfrac{y^2}{3^2}-\dfrac{z^2}{2^2}=x^2-1"

Axis X: "\dfrac{x^2}{6^2}+\dfrac{0^2}{3^2}-\dfrac{0^2}{2^2}=x^2-1"

"x^2(1- \dfrac{1}{6^2})-1=0"

"x_1=- \dfrac{6}{\sqrt35}"

"x_2=\dfrac{6}{\sqrt35}"

Axis Y:"\dfrac{0}{6^2}+\dfrac{y^2}{3^2}-\dfrac{0^2}{2^2}=0^2-1"

"\dfrac{y^2}{3^2}=-1"

So this conicoid hasn't any common points with axis Y

Axis Z:"-\dfrac{z^2}{2^2}=0^2-1"

"\dfrac{z^2}{2^2}=1"

"z^2=2^2"

"z_1=2"

"z_2=-2"